As with most bonds, consider a bond with a face value of $1,000. The bond's maturity is 6 years, the coupon rate is 7% paid semiannually, and the discount rate is 18%.

What is the estimated value of this bond today?
Answer #1

Information provided:

Face value= future value= $1,000

Time= 6 years*2= 12 semi-annual periods

Coupon rate= 7%/2= 3.5%

Coupon payment= 0.035*1,000= $35 per semi-annual period

Yield to maturity= 18%/2= 9% per semi-annual period

The value of the bond is calculated by computing the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

PMT= 35

I/Y= 9

N= 12

Press the CPT key and PV to compute the present value.

The value obtained is 606.16.

Therefore, the value of the bond today is $606.16.

In case of any query, kindly comment on the solution.